Chapter 7 Review Page 388
25.A furniture finish consists of turpentine and linseed oil. It contains twice as much turpentine as linseed oil. If you plan to make 16 fluid ounces of furniture finish, how much turpentine do you need?
Solve with Substitution, Elminination, or Graphing.

Answer by ptaylor(2048)

Let x=amount of linseed oil
And let y=amount of turpentine
Now we are told that y=2x---------------------------eq1
We are also told that:
x+y=16-------------------------------------eq2
Substitute y=2x from eq1 into eq2
x+2x=16
3x=16 divide both sides by 3
x=5 1/3 fl oz----------------------------amount of linseed oil needed
substitute this into eq1:
(5 1/3)+y=16 or
(16/3)+y=16 multiply each term by 3
16+3y=48 subtract 16 from each side
16-16+3y=48-16 collect like terms
3y=32 divide both sides by 2
y=10 2/3 fl oz------------------------------------amount of turpentine needed
CK
5 1/3 + 10/ 2/3=16
16=16
Also
2*5 1/3=10 2/3
10 2/3= 10 2/3


Hope this helps ---ptaylor